OpenStudy (anonymous):

Take the derivative y=e^(3x) *(3sinx-cosx)

OpenStudy (anonymous):

do I use the chain rule or product rule?

OpenStudy (akashdeepdeb):

Both. :D

OpenStudy (akashdeepdeb):

First use product rule. And then use chain rule.

OpenStudy (anonymous):

oh my >_<

OpenStudy (anonymous):

ok

OpenStudy (akashdeepdeb):

It is not very difficult. :)

OpenStudy (anonymous):

product rule is sufficient..

hartnn (hartnn):

chain rule just for differentiating e^3x

OpenStudy (akashdeepdeb):

You have to use chain rule for e^(3x), differentiate it wrt 3x and then multiply it with 3. That's it.

Miracrown (miracrown):

I agree with @AkashdeepDeb use both! we need chain rule for the e^(3x) and product rule for the other parts

OpenStudy (anonymous):

f(x)=e^3x f'(x)=(e^3x)' g(x)=3sinx-cosx g'(x)=3cosx+sinx ?

Miracrown (miracrown):

first u can distribute the e^3x so it splits into two separate derivatives and then you can do them one at a time and just add them together at the end

OpenStudy (anonymous):

I've been doing too much math in a row lol

Miracrown (miracrown):

and I've been helping too much in a row lol

hartnn (hartnn):

both of those are good things! :)

Miracrown (miracrown):

anyways math is never too much, too much math is always too less.

OpenStudy (anonymous):

wait so e^3x using chain rule

Miracrown (miracrown):

yes thats correct you just need to combine them

OpenStudy (anonymous):

my e^u is f and 3x is my g?

Miracrown (miracrown):

well its really just the derivative of e^3x=3e^3x

Miracrown (miracrown):

technically its the chain rule

Miracrown (miracrown):

i believe so yes

OpenStudy (anonymous):

ok, got the same result

OpenStudy (anonymous):

so then I

OpenStudy (anonymous):

(3e^3x)(3sinx-cosx)+(e^3x)(3cosx+sinx)

OpenStudy (anonymous):

is this the next step?

OpenStudy (anonymous):

@Miracrown @hartnn

hartnn (hartnn):

yes, thats correct :)

hartnn (hartnn):

factor out e^(3x) to simplify

OpenStudy (anonymous):

great. Should/can I simplify this?

hartnn (hartnn):

yes

hartnn (hartnn):

e^3x [ (3)(3sinx-cosx)+(3cosx+sinx) ]

OpenStudy (anonymous):

ah, understood. Thanks again :)

hartnn (hartnn):

that can be further simplified...i hope you do that :) and welcome ^_^

hartnn (hartnn):

infact cos x term gets cancelled!

OpenStudy (anonymous):

ok, I'm working it out

OpenStudy (anonymous):

e^3x (10sinx)

hartnn (hartnn):

absolutely correct! :)